OpenStudy (anonymous):

r(b-2)+t(b-2)= (r+t)(b-2)?

OpenStudy (anonymous):

true.. . r(b-2)+t(b-2) = (r+t)(b-2) r+t (b-2) = r+r (b-2)

OpenStudy (anonymous):

\[r(b-2)+t(b-2)=(b-2) (r+t) \]

OpenStudy (anonymous):

The b-2 factors out.

OpenStudy (anonymous):

true